OpenStudy (anonymous):

What is the solution to the equation log 2 x - log 2 4 = 2?

OpenStudy (anonymous):

using the properties of log to get the left side into a fraction

sam (.sam.):

\[\log_2x-\log_24=2?\]

OpenStudy (anonymous):

oh same base nvm no need for that

sam (.sam.):

\[\begin{array}{l}\text{Properties of log} \frac{\log (x)}{\log (2)}-2=2 \\ \text{Add }2\text{ to both sides} \\ \frac{\log (x)}{\log (2)}=4 \\ \text{Multiply both sides by }\frac{1}{\log (2)} \\ \log (x)=4 \log (2) \\ \text{Cancel logarithms} \\ x=16 \\\end{array}\]

OpenStudy (anonymous):

first convert to single log such as log (2x/24)=2 then convert to exponential form the base of log is 10 so 10^2=(2x/24) then simplify (2x/24) which is 1/12 noq you have 1x/12=100 so multiply by 12 on both side you have x=1200